OpenStudy (anonymous):

Solve the following Quadratic Equations. x2 = 49

OpenStudy (anonymous):

x^2-49 = 0

OpenStudy (callisto):

Take square root for both sides, what do you get?

OpenStudy (anonymous):

@broderick365 try it

OpenStudy (anonymous):

x = -7 x = 7

OpenStudy (anonymous):

(X-7) (X+7)

OpenStudy (callisto):

@broderick365 Actually, you don't have to factor the expression in this case. As I've mentioned, take square root for both sides: \[\sqrt{x^2} =\pm\sqrt{ 49}\] Simplify it and get your answer!

OpenStudy (anonymous):

7

OpenStudy (anonymous):

?????????

OpenStudy (callisto):

That's one of the answers. Note that there is an \(\pm\) for the right.

OpenStudy (anonymous):

+PLUS OR MINUS 7

OpenStudy (callisto):

Yes. That means x=-7 or x=7 can be the answer to the question. Got it?